Calculate the molecules/cm3 AND moles/L concentration of the following samples: Show work

a) 15 ppm O3 at 732 mmHg and 25.0 C

b) 23 ppb NO at 748 mmHg and 21.0 C
Answer #1

we will use conversions : 1 L= 103 cm3 1 atm = 760 mm of Hg and

                                    moles = mass/molecular weight , 1 mole = NA molecules or atoms, NA= 6.023 * 1023

ppm = parts per million and ppb = parts per billion

α) 15ppm : 15 in 101 gham ne, 1 Velume 7 95 L 7.95 x 1o c Ow », 0.313 mele = NA × 0.313 о-24 xlo2on.de calw, = з

b) 23 ppb 23 in 103 gham x748 - 0 94- atm 160 no. et melis el No 14+16 no et molsculs in 0-77 huslus 21 o 2.3 18.89 xIo

a) m hneees / L 0.313/795 18 99 04 mus